Let X be a random variable with probability density function fx= c1-x2,   -1<x<10,   &nb...

  1. Let X be a random variable with probability density function
    fx= c1-x2,   -1<x<10,                  otherwise4k9CeJzK0akJgHATIRzEO7jQrIUAIEAI6AmQXkz4
    1. What is the support of X?
    2. What is the value of c?
    3. Sketch the probability density function of X.
    4. Find P(X<0).
    5. Find P(X<0.5).
    6. Find P(X<2).
    7. Determine the expected value of X.
Answer #1

Let x be Probability density function, f(x) = 5 CC 1-x2) , -1<x< 1 0 0.w X is lies between -l to I © 5 668) dx = 1 & Sc(1-x²)f(x) 075 p05+ to-ois o ois i PLX PLX20 ) = $(1-x2) dx C1-82) dx 3 [(0+1).(0+!)] = 3 /1-1 - 3 x 2 0.5 | © PLX< 05]: (LY) 3 -1PLX <2) we can not calculate probability because limit of x this is (-1 toll o Expected value EXE S xf(x) dx 3 X (1-x2) dx 38